Course Name: Financial Derivatives (F1N429) Quiz Number: 1 (Lecture W 5 Chapter 1 of Pundamentals of Putures and Options Markets Date: Jan 30, 2019 Session: 12:30-12:55 PM Time Allowed: 25Mi & OL ection: FI 429 ZAYED UNIVERSITY Student details 20 Name: Please solve all questions. Question # 1 (Max. Marks-4-2+2) Define a forward contract. How is it different from futures contract? Question # 2 (Max. Marks-6-2+2+2) Suppose that price of a share of City Bank on January 31, 2007 is $20. The same day Gulf Investments Company entered into a contract with Bear and Stern Bank to buy 10,000 shares of City Bank on June 30, 2007 at the price of $22 per share. Lets suppose that price of City Bank share is $21.5/share on June 30, 2007 1. Who will loose, long or short position holder, on june 30, 2007 and how much? 2. What is the delivery/exercise price in the above-mentioned contract? 3. What will be the terminal pay off for short position holder if price of City Bank is 19/share or 21/share or 23/share?
Answer #1

Answer to question 1

A forward contract is a derivative contract by which one party agrees to buy and another party agrees to sell something on a pre-determined day at a pre-determined price in future.

It is different from a futures contract as it is not regulated through a clearing corporation and and the terms are much more flexible than future contracts.

Answer to question 2

(1) The long position will loose as it entered into a contract to buy at 22 but it is available in the market for just 21.50

(2) The exercise price is $22 as this was the agreed upon price between the two parties.

(3) At 19, the short position holder will gain (22-19) = $3 per share

At 21, the short position holder will gain (22-21) = $1 per share

At 23, the short position holder will lose (23-22) = $1 per share [Short position holder has a right to sell at 22, hence at any price below 22 he will gain and at any price above 22 he will lose]